OpenStudy (anonymous):

The gravitational attraction between 2 bodies varies inversely as the square of the distance between them. If the force of attraction is 64 pounds when the distance between the bodies is 9ft what is the distance between them when the force is 36 pounds?

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (cwrw238):

inverse variation y = k/x^2 where k is some constant so let y = the force and x = distance and plug in the given values to find k now replace k by this value , plug in the force = 36 and solve for x

OpenStudy (anonymous):

@cwrw238 36=k/x^2 would K be any number?

OpenStudy (cwrw238):

as i said you can find k by putting y = 64 and x = 9 in the formula and solving for k

OpenStudy (cwrw238):

64 = k / (9^2)

OpenStudy (cwrw238):